# Hopping Too Late: Exploring the Limitations of Large Language Models on Multi-Hop Queries

### v1

@@ Abstract
Large language models (LLMs) can solve complex multi-step problems, but little is known about how these computations are implemented internally. Motivated by this, we study how LLMs answer multi-hop queries such as “The spouse of the performer of Imagine is”. These queries require two information extraction steps: a latent one for resolving the first hop (“the performer of Imagine”) into the bridge entity (John Lennon), and one for resolving the second hop (“the spouse of John Lennon”) into the target entity (Yoko Ono). Understanding how the latent step is computed internally is key to understanding the overall computation. By carefully analyzing the internal computations of transformer-based LLMs, we discover that the bridge entity is resolved in the early layers of the model. Then, only after this resolution, the two-hop query is solved in the later layers. Because the second hop commences in later layers, there could be cases where these layers no longer encode the necessary knowledge for correctly predicting the answer. Motivated by this, we propose a novel “back-patching” analysis method whereby a hidden representation from a later layer is patched back to an earlier layer. We find that in up to 57% of previously incorrect cases there exists a back-patch that results in the correct generation of the answer, showing that the later layers indeed sometimes lack the needed functionality. Overall our methods and findings open further opportunities for understanding and improving latent reasoning in transformer-based LLMs.
@@ 1 Introduction
Using this approach, in §3 we find that the outcome of the first hop is clearly evident in the hidden representation of the end-token of the first hop. We further find that this resolving happens during the early layers, as depicted in Figure 1 (stage 1). Given the localization of the first hop, in §4 we then ask where the second hop is resolved. We find that it typically appears in the last token of the prompt only in layers after the resolution of the first hop, as further illustrated in Figure 1 (stage 3).
These observations suggest a mechanism where the first hop is resolved with the answer, that then propagates to the last token to resolve the second hop (this is depicted by stage 2 in Figure 1). Moreover, they suggest why this process may fail — If the two resolutions are done by different groups of layers, it must be that these two groups are both able to perform knowledge extraction, be it via their attention or MLP sublayers. However, since transformers have a limited amount of layers, and their functionality is different (Meng et al., 2022; Geva et al., 2023), it is quite likely that the layer at which the second hop is performed will not have the desired functionality. To verify this hypothesis, in §5, we propose an analysis method named back-patching. The crux of the method involves taking a hidden representation from a later layer, patching it back into the same position of the same prompt in an earlier layer, and then continuing the forward pass. In a way, this allows the model more layers to finish the computation without the need for additional parameters or training (Petty et al., 2024). Unlike previous methods (Sakarvadia et al., 2023; Li et al., 2024a), back-patching makes no assumption on the structure of the prompt and does not require any knowledge of the participating entities. Testing back-patching on queries initially predicted incorrectly, we find that up to 57% of incorrect cases have the potential to be correctly predicted using back-patching when choosing the optimal source and target layers.
@@ 2.2 Dataset
We are specifically interested in understanding the differences between when the model completes the two-hop query correctly and incorrectly. Therefore we generate two data subsets for these cases accordingly.
@@
The second subset includes cases where the model correctly answers both the first and second hop in isolation, but fails to answer the full two-hop query. As an example for the filtering procedure, given the same query above, we verify that the model correctly predicts John Lennon for the first hop “The performer of Imagine is” and Yoko Ono for the second hop “The spouse of John Lennon is”, but fails to predict Yoko Ono for the two-hop query “The spouse of the performer of Imagine is”. We focus on these cases, because we are interested to understand why answering the two-hop query fails despite the model “knowing” the two separate facts. As we show later, this failure is likely due to the first fact being resolved too late. In a similar fashion to the first setting, we then randomly sample 50 examples of each bridge entity type (in order to keep the correct and incorrect subsets of approximately the same size).
Table 1 reports the amount of examples used in experiments by model and settings.
@@ 2.3 Models
We analyze the following LLaMA models: LLaMA 2 7B and 13B (Touvron et al., 2023) and LLaMA 3 8B and 70B (Meta AI, 2024). The 7B and 8B models have 32 layers, while the 13B model has 40 layers and the 70B model has 80 layers.
@@ 3 Localizing First Hop Resolution
One could imagine multiple possible ways for a model to solve multi-hop queries. One such way would be to effectively treat the two-hop query as a single relation, enabling the extraction of the final answer without the need to recall $e_{2}$. Despite this, if the model truly is performing latent reasoning, the most intuitive way to do so would start by resolving the first hop at some point within the computation. In order to localize where this happens, we use the Patchscopes analysis method. Our results show that the model indeed resolves the first hop during the early layers, at the position of $t_{1}$.
Additionally, as the model’s prediction is emitted from the last position of the prompt, critical information regarding the now-resolved bridge entity must propagate to this token. We further explore and validate this process in Appendix B.
3.1 Interpreting Hidden Representations Using Patchscopes
@@ 3.3 Results
Model Setting $t_{1}$ encodes $e_{2}$ $t_{2}$ encodes $e_{3}$
Cases Layer Cases Layer
LLaMA 2 7B Correct $56.1\%$ $8.9$ $73.4\%$ $16.2$
Incorrect $41.8\%$ $9.1$ $47.0\%$ $17.5$
LLaMA 2 13B Correct $49.2\%$ $7.7$ $71.1\%$ $16.9$
Incorrect $48.6\%$ $7.4$ $33.4\%$ $16.9$
LLaMA 3 8B Correct $46.4\%$ $8.9$ $74.4\%$ $13.5$
Incorrect $46.9\%$ $11.9$ $40.4\%$ $17.2$
LLaMA 3 70B Correct $63.5\%$ $27.1$ $86.7\%$ $30.7$
Incorrect $57.8\%$ $29.2$ $50.0\%$ $35.9$
Table 2: The results of Patchscopes executions. The table reports the percentage of cases where a target entity was successfully decoded from a source position. The table additionally reports the mean of the first layer where the target entity was successfully decoded.
@@
Figure 3: Heat-map of the layers where the entity description task successfully decodes $e_{2}$ from the position of $t_{1}$. The percentages are out of all successful decodings for LLaMA 2 13B run on correctly answered cases.
Table 2 present the percentage of cases $e_{2}$ was observed in the hidden representation of $t_{1}$. We refer to these as cases where $e_{2}$ was successfully resolved. We note the non-trivial percentages of 41%-63% across all models and settings. Additionally, we note the general drop in resolved cases when the model fails to correctly complete the two-hop query, but that this setting still displays a surprisingly high success rate. Next, we examine the mean layer at which $e_{2}$ first appears in both correct and incorrect cases. We find that in the incorrect cases the resolving generally occurs at higher layers compared to those of the correct cases. This difference alludes to how early the resolving happens potentially playing a part in successfully answering two-hop queries. We further verify this using back-patching in §5.
Figure 2 displays the proportion of cases per layer where $e_{2}$ is first successfully decoded from $t_{1}$ using Patchscopes. We observe that the resolving occurs almost exclusively during the early layers. Interestingly, this implies that the knowledge required for this resolution must reside in these earlier layers. Similar figures for additional models are provided in Appendix C. The plot additionally shows the curve regarding the resolving of $e_{3}$ from $t_{2}$, which is discussed in §4.
This finding can further be seen in Figure 3, which displays a heat-map of the layers where Patchscopes successfully decodes $e_{2}$ from $t_{1}$. Two hot-spots are apparent. The first corresponds to success cases with an early source layer, showing that $e_{2}$ generally appears at this position during the early layers. The second corresponds to later source layers that display success almost exclusively when patched after the final layer, which is roughly equivalent to vocabulary projection (nostalgebraist, 2020). This is a clear advantage of Patchscopes, as despite the entity not being among the top tokens in the projection during early layers, it is in fact encoded in the hidden representation. See Appendix D for heat-maps of all models.
@@ 4 Second Hop is Resolved at Last Position
In §3, we observe that the first hop is often resolved in position $t_{1}$. Intuitively, we’d expect the computation to use this result towards resolving the second hop (i.e., the final answer). In this section, we study this by finding where $e_{3}$ first emerges and how it relates to the first emergence of $e_{2}$. We find that $e_{3}$ indeed emerges predominantly after $e_{2}$ at the position of $t_{2}$. Furthermore, we search for the final piece of the pathway, the predicted token. By projecting the residual updates made by the attention and MLP sublayers on the hidden representation of $t_{2}$, we find that the MLP sublayers play the larger part in conjuring the predicted token.
@@ 4.2 Experiment
We first employ Patchscopes on $t_{2}$, this time checking if the target entity $e_{3}$ was encoded at this position. Additionally, we project the sublayer updates performed on $t_{2}$, and check if the token with the highest probability in the projection is the first token of the generated text. If so, we conclude that the update of a particular sublayer promotes the final prediction.
@@ 4.3 Results
Model Setting Attention MLP
@@
Table 2 contains the percentage of cases $e_{3}$ was decoded from $t_{2}$, effectively resolving the second hop. As expected, these percentages are high for cases the model predicted correctly (71%-86%) [^3] We attribute the gap from 100% to the approximations involved in decoding a hidden representation. and low for cases the model predicted incorrectly (33%-50%). Additionally, it is apparent from both Table 2 and Figure 2 that the target entity is resolved by the mid-upper layers at the position of $t_{2}$. This suggests that the information required to perform the second hop exists in the parameters of these layers.
Inspecting the sublayer projections in Table 3, we find that the MLP sublayers play a larger role in promoting the predicted token compared to that of the attention sublayers. Despite this, the non negligible percentages achieved by the attention sublayers could suggest multiple pathways, in agreement with previous work (Geva et al., 2023; Merullo et al., 2023) that considered both possibilities. Looking at the mean layers at which the prediction is first promoted, one can see that these promotions occur in the upper layers, after the last token has been resolved to $e_{3}$.
@@
After describing all major pathway stages, we turn to a comparison between correct and incorrect cases and observe the following pattern. Generally, in incorrect cases, the entities seem to be resolved later while the information seems to propagate and be extracted earlier. Figure 4 displays this pattern in LLaMA 3 8B, plotting the first observed layers of each stage (we refer to Appendix E for plots of additional models). This gives evidence of the importance of how early the first hop is resolved.
@@ 5 Back-patching Improves Two-Hop Performance
Our results consistently display a sequential nature of computation, where the first hop is resolved into the bridge entity by the lower layers, which is then used to answer the second hop in the upper layers. These results suggest that when answering two-hop queries fail, it could be because the first-hop is resolved at layers that no longer contain the information needed to resolve the second hop. In order to verify this hypothesis, we test the following intriguing conjecture. We know that the earlier layers certainly contain knowledge, since they are regularly used to resolve the first hop. Then, we ask what will happen if we take a hidden representation from a later layer and “feed” it into these earlier layers. Will the correct information now be extracted? As our results next show, this is often the case. This “back-patching” analysis thus provides strong support that the success in two hops requires the later layers to have the same functionality as earlier ones.
@@ 5.2 Experiment
We perform the back-patching experiment twice, once on the token $t_{1}$ and once on the token $t_{2}$. In order to quantify the potential of back-patching, we check whether one of source-target pairs resulted in the generation of the correct answer. We then perform an analysis of the successful source and target layers.
@@ 5.3 Results
Table 4 presents the back-patching results. First we note the high success rates, correctly answering between 32%-57% of incorrect cases while keeping the correct cases at 100% across models and token positions. This means that in many cases rerunning specific parts of the computation is enough in order to successfully extract the target information. On a more practical note, this also shows that significant gains can be made by correctly choosing the source and target layers, which we leave for future work.
One can additionally observe that back-patching $t_{2}$ achieves a higher success rate than back-patching $t_{1}$, which may point at the stages related to the last position of the prompt as being the more dominant failure point in answering two-hop queries. Moreover, the fact that back-patching $t_{2}$ does not assume any information about the structure of the prompt shows an additional strength of the method.
Figure 5(a) displays the heat-map of the layers where back-patching $t_{1}$ succeeds for LLaMA 2 13B (see Appendix F for additional models). We first note the higher success rate when the source layers are the model’s lower layers. These lower layers correlate to the resolving of the first hop, therefore back-patching them allows the model a better chance at solving the first hop. Second, we note the non existent success rate of the higher target layers. Unsurprisingly, in this case the model does not gain enough time to correctly complete the pathway, resulting in a low success rate.
Figure 5(b) displays the heat-map of the layers where back-patching $t_{2}$ succeeds for LLaMA 2 13B. Unlike when back-patching $t_{1}$, in this case the higher success rate can be found in the the upper source layers. These layers correlate to the resolving of the second hop at the last token position, hence back-patching the hidden representation taken from them allows the model a better chance at solving the second hop. Furthermore, we note the higher success rate in lower target layers. This allows the model to gain time to resolve the second hop. This also gives the model access to parametric knowledge that might only be encoded in the lower layers, allowing the model to fully utilize its whole knowledge-base.

### v2

@@ Abstract
Large language models (LLMs) can solve complex multi-step problems, but little is known about how these computations are implemented internally. Motivated by this, we study how LLMs answer multi-hop queries such as “The spouse of the performer of Imagine is”. These queries require two information extraction steps: a latent one for resolving the first hop (“the performer of Imagine”) into the bridge entity (John Lennon), and another for resolving the second hop (“the spouse of John Lennon”) into the target entity (Yoko Ono). Understanding how the latent step is computed internally is key to understanding the overall computation. By carefully analyzing the internal computations of transformer-based LLMs, we discover that the bridge entity is resolved in the early layers of the model. Then, only after this resolution, the two-hop query is solved in the later layers. Because the second hop commences in later layers, there could be cases where these layers no longer encode the necessary knowledge for correctly predicting the answer. Motivated by this, we propose a novel “back-patching” analysis method whereby a hidden representation from a later layer is patched back to an earlier layer. We find that in up to 66% of previously incorrect cases there exists a back-patch that results in the correct generation of the answer, showing that the later layers indeed sometimes lack the needed functionality. Overall, our methods and findings open further opportunities for understanding and improving latent reasoning in transformer-based LLMs.
@@ 1 Introduction
Using this approach, in §3 we find that the outcome of the first hop is clearly evident in the hidden representation of the end-token of the first hop. We further find that resolving the first hop happens during the early layers, as depicted in Figure 1 (stage 1). Given the localization of the first hop, in §4 we then ask where the second hop is resolved. We find that it typically appears in the last token of the prompt only in layers after the resolution of the first hop, as further illustrated in Figure 1 (stage 3).
These observations suggest a mechanism where the first hop is resolved with the answer, that then propagates to the last token to resolve the second hop (as depicted by stage 2 in Figure 1). We explore this propagation in §5. Moreover, they suggest why this process may fail — If the two resolutions are done by different groups of layers, it must be that these two groups are both able to perform knowledge extraction, be it via their attention or MLP sublayers. However, since transformers have a limited amount of layers, and their functionality is different (Meng et al., 2022; Geva et al., 2023), it is quite likely that the layer at which the second hop is performed will not have the desired functionality. To verify this hypothesis, in §6, we propose an analysis method named back-patching. The crux of the method involves taking a hidden representation from a later layer, patching it back into the same position of the same prompt in an earlier layer, and then continuing the forward pass. In a way, this allows the model more layers to finish the computation without the need for additional parameters or training (Petty et al., 2024). Unlike previous methods (Sakarvadia et al., 2023; Li et al., 2024a), back-patching makes no assumption on the structure of the prompt and does not require any knowledge of the participating entities. Testing back-patching on queries initially predicted incorrectly, we find that up to 66% of incorrect cases have the potential to be correctly predicted using back-patching when choosing the optimal source and target layers.
@@ 2.2 Dataset
We are specifically interested in understanding the differences between cases where the model completes the two-hop query correctly and incorrectly. Therefore we generate two dataset subsets per model for these cases accordingly.
@@
The second subset includes cases where the model correctly answers both the first and second hop in isolation, but fails to answer the full two-hop query. As an example for the filtering procedure, given the same query above, we verify that the model correctly predicts John Lennon for the first hop “The performer of Imagine is” and Yoko Ono for the second hop “The spouse of John Lennon is”, but fails to predict Yoko Ono for the two-hop query “The spouse of the performer of Imagine is”. We focus on these cases, because we are interested in understanding why answering the two-hop query fails despite the model “knowing” the two separate facts. As we show later, this failure is likely due to the first fact being resolved too late. In a similar fashion to the first setting, we then randomly sample 50 examples of each bridge entity type (in order to keep the correct and incorrect subsets of approximately the same size).
Table 1 reports the amount of examples used in experiments by model and setting.
@@ 2.3 Models
We analyze the following models: LLaMA 2 7B and 13B (Touvron et al., 2023), LLaMA 3 8B and 70B (Meta AI, 2024), and Pythia 6.9B and 12B (Biderman et al., 2023). The 6.9B, 7B and 8B models have 32 layers, the 12B model has 36 layers, the 13B model has 40 layers, and the 70B model has 80 layers.
@@ 3 Localizing First Hop Resolution
One could imagine several possible ways for a model to solve multi-hop queries. One such way would be to effectively treat the two-hop query as a single relation, enabling the extraction of the final answer without the need to recall $e_{2}$. Despite this, if the model truly performs latent reasoning, the most intuitive way to do so would start by resolving the first hop at some point within the computation. In order to localize where this happens, we use the Patchscopes analysis method. Our results show that the model indeed resolves the first hop during the early layers, at the position of $t_{1}$.
3.1 Interpreting Hidden Representations
@@ 3.3 Results
Model Subset $e_{2}$ from $t_{1}$ $e_{2}$ from $t_{2}$ $e_{3}$ from $t_{2}$
Cases Layer Cases Layer Cases Layer
LLaMA 2 Corr. $56\%$ $8.9$ $41\%$ $15.3$ $73\%$ $16.2$
7B Incorr. $41\%$ $9.1$ $36\%$ $16.4$ $47\%$ $17.5$
LLaMA 2 Corr. $49\%$ $7.7$ $34\%$ $18.2$ $71\%$ $16.9$
13B Incorr. $48\%$ $7.4$ $37\%$ $17.9$ $33\%$ $16.9$
LLaMA 3 Corr. $46\%$ $8.9$ $37\%$ $14.3$ $74\%$ $13.5$
8B Incorr. $46\%$ $11.9$ $25\%$ $15.3$ $40\%$ $17.2$
LLaMA 3 Corr. $63\%$ $27.1$ $46\%$ $35.2$ $86\%$ $30.7$
70B Incorr. $57\%$ $29.2$ $46\%$ $34.5$ $50\%$ $35.9$
Pythia Corr. $75\%$ $5.4$ $75\%$ $11.4$ $80\%$ $14.1$
6.9B Incorr. $78\%$ $4.9$ $67\%$ $9.2$ $65\%$ $13.6$
Pythia Corr. $73\%$ $5.0$ $66\%$ $12.2$ $77\%$ $13.5$
12B Incorr. $61\%$ $6.3$ $42\%$ $12.9$ $52\%$ $17.2$
Table 2: The results of Patchscopes executions. The table reports the percentage of cases where a target entity was successfully decoded from a source position. The table additionally reports the mean of the layer where the target entity was first successfully decoded.
@@
Figure 3: Heat-map of the layers where Patchscopes successfully decodes $e_{2}$ from the position of $t_{1}$. The percentages are out of all successful decodings for LLaMA 2 13B run on correctly answered cases.
Table 2 presents the percentage of cases where $e_{2}$ was observed in the hidden representation of $t_{1}$. We refer to these as cases where $e_{2}$ was successfully resolved. We note the non-trivial percentages of 41%-78% across all models and settings. Additionally, we note the general drop in resolved cases when the model fails to correctly complete the two-hop query, but that this setting still displays a surprisingly high success rate. Next, we examine the mean layer at which $e_{2}$ first appears in both correct and incorrect cases. We find that in the incorrect cases the resolving generally occurs at higher layers than those of the correct cases. This difference alludes to how early the resolving happens potentially playing a part in successfully answering two-hop queries. We further verify this using the back-patching analysis method in §6.
Figure 2 displays the proportion of cases per layer where $e_{2}$ is first successfully decoded from $t_{1}$ using Patchscopes. We observe that the resolving occurs almost exclusively during the early layers. Interestingly, this implies that the knowledge required for this resolution must reside in these earlier layers. Similar figures for additional models are provided in Appendix B. The plot additionally shows the curve regarding the resolving of $e_{3}$ from $t_{2}$, which is discussed in §4.
This finding can further be seen in Figure 3, which displays a heat-map of the layers where Patchscopes successfully decodes $e_{2}$ from $t_{1}$. Two hot-spots are apparent. The first corresponds to success cases with an early source layer, showing that $e_{2}$ generally appears at this position during the early layers. The second hot-spot corresponds to later source layers that display success almost exclusively when patched after the final layer, which is roughly equivalent to vocabulary projection (nostalgebraist, 2020). This is a clear advantage of Patchscopes, as despite the entity not being among the top tokens in the projection during early layers, it is in fact encoded in the hidden representation. See Appendix C for heat-maps of all models.
@@ 4 Second Hop is Resolved at Last Position
In §3, we observe that the first hop is often resolved in position $t_{1}$. Intuitively, one would expect the computation to use this result towards resolving the second hop (i.e., the final answer). In this section, we study this proposition by finding where $e_{3}$ first emerges and how it relates to the first emergence of $e_{2}$. We find that $e_{3}$ indeed emerges predominantly after $e_{2}$ at the position of $t_{2}$. Furthermore, we search for the final part of the pathway, the predicted token. By projecting the residual updates made by the attention and MLP sublayers on the hidden representation of $t_{2}$, we find that the MLP sublayers play the larger part in conjuring the predicted token.
@@ 4.2 Experiment
We first employ Patchscopes on $t_{2}$, this time checking if the target entity $e_{3}$ is encoded at this position. Additionally, we project the sublayer updates performed on $t_{2}$, and check if the token with the highest probability in the projection is the first token of the generated text. If so, we conclude that the update of a particular sublayer promotes the final prediction.
@@ 4.3 Results
Model Subset Attention MLP
@@
Pythia 6.9B Correct $11.5\%$ $21.8$ $36.4\%$ $20.5$
Incorrect $28.2\%$ $23.4$ $33.6\%$ $21.2$
Pythia 12B Correct $27.9\%$ $22.2$ $38.9\%$ $22.4$
Incorrect $43.2\%$ $23.6$ $47.0\%$ $23.3$
@@
Table 2 contains the percentage of cases $e_{3}$ was decoded from $t_{2}$, effectively resolving the second hop. As expected, these percentages are high for cases in which the model predicted correctly (71%-86%) [^3] We attribute the gap from 100% to the approximations involved in decoding a hidden representation. and low for cases the model predicted incorrectly (33%-65%). Additionally, it is apparent from both Table 2 and Figure 2 that the target entity is resolved by the mid-upper layers at the position of $t_{2}$. This suggests that the information required to perform the second hop exists in the parameters of these layers.
Inspecting the sublayer projections in Table 3, we find that the MLP sublayers play a larger role in promoting the predicted token compared to that of the attention sublayers. Despite this, the non negligible percentages achieved by the attention sublayers could suggest multiple pathways, in agreement with previous work (Geva et al., 2023; Merullo et al., 2024) that considered both possibilities. Looking at the mean layers at which the prediction is first promoted, one can see that these promotions occur in the upper layers, after the last token has been resolved to $e_{3}$.
5 Information Propagates to Last Token
5.1 Method
In similar fashion to previous work (Wang et al., 2023b; Geva et al., 2023), we block hidden representations from attending to other representations at specific layers and test if the model’s final generation changes. We perform the knockout by setting the mask of the attention from a source to a target position at a specific layer to $\infty$- ∞, causing this attention edge to contribute nothing to the residual update performed on the source representation. We experiment with blocking a window of layers as previous work (Geva et al., 2023) shows that information propagation is spread among multiple layers.
5.2 Experiment
We run attention knockout blocking information originating at $t_{1}$ from propagating to $t_{2}$. We block a window of 7 layers in order to capture information flowing through multiple layers. If blocking a set of layers causes the model to no longer correctly predict the answer (in the incorrect setting we check for a change in the generated text), we consider these layers to contain information critical to the prediction.
5.3 Results
Model Subset Detected Mean Layer
LLaMA 2 7B Correct $85.82\%$ $12.83$
Incorrect $95.44\%$ $8.10$
LLaMA 2 13B Correct $81.94\%$ $15.45$
Incorrect $71.11\%$ $16.99$
LLaMA 3 8B Correct $78.36\%$ $10.15$
Incorrect $82.74\%$ $7.45$
LLaMA 3 70B Correct $90.39\%$ $29.14$
Incorrect $94.28\%$ $20.06$
Pythia 6.9B Correct $96.53\%$ $11.94$
Incorrect $95.04\%$ $8.06$
Pythia 12B Correct $94.18\%$ $11.23$
Incorrect $93.81\%$ $9.28$
Table 4: Percentages and mean layers of cases where critical information was first detected propagating from $t_{1}$ to $t_{2}$ using attention knockout, projection, or Patchscopes.
Table 2 contains the percentage of cases $e_{2}$ was decoded from $t_{2}$, showing that this position indeed represents the bridge entity in 25%-75% of all cases. Comparing between the results of Patchscopes on $t_{2}$ to those on $t_{1}$, we observe a stark difference in the successful source layers. The results show that the $t_{2}$ first encodes $e_{2}$ during the mid-upper layers, while $t_{1}$ does so in the lower layers. This can further be seen in the Patchscopes heat-maps in Appendix C. This gives additional reason to believe that a sequential computation is taking place.
@@ 5.3 Results
After describing all major pathway stages, we turn to a comparison between the correct and incorrect cases and observe the following pattern. Generally, in incorrect cases, the entities seem to be resolved later while the information seems to propagate and be extracted earlier. Figure 4 displays this pattern in LLaMA 3 8B, plotting the first observed layers of each stage (we refer to Appendix D for plots of additional models). This gives evidence of the importance of how early the first hop is resolved.
@@ 6 Back-patching Improves Two-Hop Performance
Our results consistently display a sequential nature of computation, where the first hop is resolved into the bridge entity by the lower layers, which is then used to answer the second hop in the upper layers. These results suggest that when answering two-hop queries fail, it could be because the first-hop is resolved at layers that no longer contain the information needed to resolve the second hop. In order to verify this hypothesis, we test the following intriguing conjecture. We know that the earlier layers certainly contain knowledge, since they are regularly used to resolve the first hop. Then, we ask what will happen if we take a hidden representation from a later layer and “feed” it into these earlier layers. Will the correct information now be extracted? As our results next show, this is often the case. This “back-patching” analysis thus provides strong support that the success in two hop queries require the later layers to have the same functionality as earlier ones.
@@ 6.2 Experiment
We perform the back-patching experiment twice, once on the token $t_{1}$ and once on the token $t_{2}$. In order to quantify the potential of back-patching, we check whether one of the source-target pairs resulted in the generation of the correct answer. We then perform an analysis of the successful source and target layers.
@@ 6.3 Results
Pythia 6.9B Correct $100\%$ $100\%$
Incorrect $66.33\%$ $56.43\%$
Pythia 12B Correct $100\%$ $100\%$
Incorrect $63.17\%$ $61.82\%$
@@
Table 5 presents the back-patching results. First we note the high success rates, correctly answering between 32%-66% of incorrect cases while keeping the correct cases at 100% across models and token positions. This means that in many cases rerunning specific parts of the computation is enough in order to successfully extract the target information. On a more practical note, this also shows that significant gains can be made by correctly choosing the source and target layers, which we leave for future work.
One can additionally observe that in LLaMA models back-patching $t_{2}$ achieves a higher success rate than back-patching $t_{1}$, which may point at the stages related to the last position of the prompt as being the more dominant failure point in answering two-hop queries. Moreover, this higher success rate is achieved despite the fact that back-patching $t_{2}$ does not assume any information about the structure of the prompt. We believe that this displays an additional strength of the method.
Figure 5(a) displays the heat-map of the layers where back-patching $t_{1}$ succeeds for LLaMA 2 13B (see Appendix E for additional models). We first note the higher success rate when the source layers are the model’s lower layers. These lower layers correlate to the resolving of the first hop, therefore back-patching them allows the model a better chance at solving the first hop. Second, we note the non existent success rate of the higher target layers. Unsurprisingly, in this case the model does not gain enough time to correctly complete the pathway, resulting in a low success rate.
Figure 5(b) displays the heat-map of the layers where back-patching $t_{2}$ succeeds for LLaMA 2 13B. Unlike when back-patching $t_{1}$, in this case the higher success rate can be found in the the upper source layers. These layers correlate to the resolving of the second hop at the last token position. Hence, back-patching a hidden representation taken from a later layer allows the model a better chance at solving the second hop. Furthermore, we note the higher success rate in lower target layers. This allows the model to gain time to resolve the second hop. This also gives the model access to parametric knowledge that might only be encoded in the lower layers, allowing the model to fully utilize its whole knowledge-base.